# Approaches to critical point theory via sequential and parametrized topological complexity

### v2

@@ 8. Equivariant fiber bundles and vertical proportionality
In contrast to the setting of sequential TCs in section 2 to 4, our Lusternik-Schnirelmann-type results for sequential parametrized setting in the last sections had the downside that some restrictive conditions have to be put on the chosen quantities, namely
@@
    Let $b_{0}$, $U$ and $\psi$ be given as in the proof of a). Then
    $f(\psi^{-1}(b,x))=f(\Phi(s(b),x))=f(x)\qquad\forall b\in U$
    by the $G$-invariance of $f$.
@@
$\Phi_{g}\colon E\to E,\quad\Phi_{g}(x):=\Phi(g,x)\quad\forall g\in G,$
$\Phi^{x}\colon G\to E,\quad\Phi^{x}(g):=\Phi(g,x)\quad\forall x\in E.$
@@
  • assume that $E$ admits and is equipped with a $\Phi$-invariant Riemannian metric, which we shall fiberwise denote by ,$_{x}$, $x\in E$,
@@
[…]
@@
(8.2) $\nabla f(\Phi_{g}(x))=(D\Phi_{g})_{x}(\nabla f(x))\quad\forall g\in G,\ x\in E.$
Since the metric on $E$ is $G$-invariant and $p$ is $G$-equivariant, the differential of $\Phi_{g}$ maps vertical tangent spaces with respect to $p$ isomorphically onto one another for each $g\in G$. One derives from this observation and from (8.2) that
@@
$\nabla f(x),v\right>=\left<\nabla f(h(x)),Dh_{x}(v)\right>\ \ \forall x\in E|_{U},\ v\in T_{x}E.$
@@
$\nabla f(x),v\right>_{x}$ =$(D\Phi_{\eta(x)})_{x}(\nabla^{\mathrm{ver}}f(x)),(D\Phi^{x})_{\eta(x)}(D\eta_{x}(v))+(D\Phi_{\eta(x)})_{x}(v)\right>_{h(x)}$
=$\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(v)+v\right>_{x},$
@@
$\|\nabla f(x)\|^{2}$ =$\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\right>_{x}+\left<\nabla^{\mathrm{ver}}f(x),\nabla f(x)\right>_{x}$
=$\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\right>_{x}+\|\nabla^{\mathrm{ver}}f(x)\|^{2},$
since $\nabla^{\mathrm{ver}}f(x),\nabla f(x)\right>=\left<\nabla^{\mathrm{ver}}f(x),\nabla^{\mathrm{ver}}f(x)\right>$, for each $x\in E|_{U}$. Using the Cauchy-Schwarz inequality and the $G$-invariance of the metric, we derive that
$\|\nabla f(x)\|^{2}$ $\leq\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}$
$\leq\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|(D\Phi^{x})_{\eta(x)}\|_{\mathrm{op}}\cdot\|D\eta_{x}(\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}.$
Since $\Phi$ is $f$-controlled with respect to $C$, we obtain
$\|\nabla f(x)\|^{2}$ $\leq C(f(x))\cdot\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|D\eta_{x}(\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}$
$\leq C(f(x))\cdot\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|D\eta_{x}\|_{\mathrm{op}}\cdot\|\nabla f(x)\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}.$
@@
Since the chosen metric on $G$ is bi-invariant, this yields that $\|D\mathrm{inv}_{g}\|_{\mathrm{op}}=1$ for all $g\in G$. We derive that
$\|D\eta_{x}\|_{\mathrm{op}}\leq\|Ds_{p(x)}\|_{\mathrm{op}}\qquad\forall x\in f^{\lambda}\cap E|_{U}.$
Inserting this into the above inequality shows the claim. ∎
@@
Let $r\in\mathbb{N}$ with $r\geq 2$. Assume that $B$ is a closed manifold and that $\varphi$ is a transitive $G$-action on $B$. Let $f\in C^{1}(E)$ be $G$-invariant and bounded from below. If $(E,\left<\cdot,\cdot\right>)$ is complete, if $(f,\left<\cdot,\cdot\right>)$ satisfies the PS condition and if $\Phi$ is $f$-controlled, then
@@
Since $E$ is a closed manifold, $(E,\left<\cdot,\cdot\right>)$ is complete and $(f,\left<\cdot,\cdot\right>)$ satisfies the PS condition. Moreover, $\Phi$ is $f$-controlled, see Remark 8.4. Since $E$ is compact, $f$ attains its maximum $\lambda$, so that $p=p|_{f^{\lambda}}$. Then claim is thus an immediate consequence of Theorem 8.7. ∎
In the rest of this section, we want to apply Corollary 8.8 to unit tangent bundles of odd-dimensional spheres $\mathbb{S}^{2n-1}\subset\mathbb{R}^{2n}$. A general upper bound for the parametrized TCs of unit sphere bundles of vector bundles has been obtained by Farber and Weinberger in [FW23] and extended to sequential TCs by Farber and Paul in [FP24]. Using our approach via critical point theory, we will be able improve on these upper bounds for unit tangent bundles of odd-dimensional spheres and compute the exact values of their sequential parametrized TCs in certain cases.
In the following we fix $n\in\mathbb{N}$ and will use the canonical identification $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$ without always spelling it out explicitly. We further view $\mathbb{S}^{2n-1}$ as equipped with the metric induced by the standard metric on $\mathbb{R}^{2n}$ and denote its unit tangent bundle by
$p\colon U\mathbb{S}^{2n-1}\to\mathbb{S}^{2n-1}.$
Explicitly,
@@
[…]
@@
(8.6) $\langle x_{1},y_{1}\rangle=\langle x_{2},y_{2}\rangle=0\quad\text{as well as }\quad\langle x_{2},y_{1}\rangle+\langle x_{1},y_{2}\rangle=0.$
@@
(8.8) $\operatorname{\mathsf{TC}}_{r}[p]\leq\sum_{\mu\in\mathbb{R}}\max\Big{\{}\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{2n-1}}\dots\times_{\mathbb{S}^{2n-1}}\times\operatorname{Crit}_{\mu_{r}}f)\ \Big{|}\ \sum_{i=1}^{r}\mu_{i}=\mu\Big{\}},$
where $\operatorname{\mathsf{TC}}_{r}[p]:=\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{2n-1}\to\mathbb{S}^{2n-1}]$. We want to further compute this upper bound on $TC_{r}[p]$ and study the critical sets of $f$ in greater detail for this purpose.
@@
To show the converse inclusion, let $X=(x_{1},x_{2})\in U\mathbb{S}^{2n-1}$ with $x_{2}\notin\{ix_{1},-ix_{1}\}$. Set $Y=(0,w)$ with $w=ix_{1}-\langle ix_{1},x_{2}\rangle x_{2}$. One verifies that indeed $Y\in T_{X}U\mathbb{S}^{2n-1}$ and that
$Df_{X}(Y)=1-\langle ix_{1},x_{2}\rangle^{2}\neq 0.$
Hence, $X\not\in\operatorname{Crit}f$, which shows that any critical point of $f$ is of the form derived above. ∎
Let $r\in\mathbb{N}$ with $r\geq 2$. We have seen in the previous Lemma that the critical values of $f$ are $1$ and $-1$, so the fiber products of the critical manifolds occurring in (8.8) are of the form
@@
We want to explicitly compute the sequential parametrized subspace complexities that occur on the right-hand side of this inequality in the case of even $n$, i.e. for spheres of dimension $14m-$ 1 for some $m\in\mathbb{N}$. Eventually, we will show the following.
@@
For each $m\in\mathbb{N}$, it holds that
@@
In particular, it holds that
(8.11) $\langle x,jx\rangle=\langle ix,jx\rangle=0\quad\text{and}\quad\|jx\|^{2}=\|x\|^{2}\qquad\forall x\in\mathbb{R}^{4m}.$
@@
Let $m\in\mathbb{N}$ and let $f\colon\mathbb{S}^{4m-1}\to\mathbb{R}$ be given as in (8.7). Let $\mu_{1},\ldots,\mu_{r}\in\{+1,-1\}$ and consider the manifold
$\mathcal{C}=\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{2}}f\times_{\mathbb{S}^{4m-1}}\ldots\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{r}}f.$
Then $\operatorname{\mathsf{TC}}_{r,p}(\mathcal{C})=1$.
@@
Consider the map $B\colon\mathbb{S}^{4m-1}\to U\mathbb{S}^{4m-1}$, $B(x)=jx$. We obtain from (8.11) that $B$ is well-defined and satisfies $\langle A(x),B(x)\rangle=0$ for all $x\in\mathbb{S}^{4m-1}$.
Let $u=((x,v_{1}),\ldots,(x,v_{r}))\in\mathcal{C}$, so that $v_{j}\in\{ix,-ix\}$ for all $j\in\{1,2,\ldots,r\}$. We define a path $\sigma_{u}\colon I\to U\mathbb{S}^{n}$ as follows. For all $j\in\{0,1,\ldots,r-2\}$ and $t\in[\frac{j}{r-1},\frac{j+1}{r-1}]$ we put
@@
We note that the unit tangent bundle $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ is fiber homotopy equivalent to a trivial fibration if and only if $m\in\{1,2\}$, so that Theorem 8.10 can not be derived by elementary arguments for $m\geq 3$.
More precisely, as shown in [JW54, Theorem 1.1], $U\mathbb{S}^{4m-1}$ has the homotopy type of $\mathbb{S}^{4m-1}\times\mathbb{S}^{4m-2}$ if and only if there exists an element of Hopf invariant one in $\pi_{8m-1}(\mathbb{S}^{4m})$. By [Ada60] this is the case only for $m\in\{1,2\}$. Thus, $p$ is not fiber homotopy trivial for $m\geq 3$.

### v4

@@ 8. Equivariant fiber bundles and vertical proportionality
In contrast to the setting of sequential TCs in section 2 to 4, our Lusternik-Schnirelmann-type results for sequential parametrized setting in the last sections had the downside that some restrictive conditions have to be assumed, namely
@@
    Let $b_{0}$, $U$ and $\psi$ be given as in the proof of a). Then $f(\psi^{-1}(b,x))=f(\Phi(s(b),x))=f(x)$ for all $b\in U$ by the $G$-invariance of $f$.
@@
$\Phi_{g}\colon E\to E,\quad\Phi_{g}(x):=\Phi(g,x)\quad\forall g\in G,\qquad\Phi^{x}\colon G\to E,\quad\Phi^{x}(g):=\Phi(g,x)\quad\forall x\in E.$
@@
  • assume that $E$ admits and is equipped with a $\Phi$-invariant Riemannian metric, which we shall fiberwise denote by $\left<\cdot,\cdot\right>_{x}$, $x\in E$,
@@
The fiberwise norms on the tangent spaces of $E$ and $G$, respectively, induced by these metrics shall both be denoted by $\|\cdot\|$. We further denote the gradient of $f\in C^{1}(E)$ with respect to $\left$$\cdot$$\cdot\right$> by $\nabla f$ and let its vertical part $\nabla^{\mathrm{ver}}f$ be defined with respect to $p$ and $\left$$\cdot$$\cdot\right$>. We further let $\|\cdot\|_{\mathrm{op}}$ denote the operator map of a linear map between normed vector spaces whenever it is apparent which normed spaces we are referring to.
@@
$\nabla f(\Phi_{g}(x))=(D\Phi_{g})_{x}(\nabla f(x))\quad\forall g\in G,\ x\in E.$
Since the metric on $E$ is $G$-invariant and $p$ is $G$-equivariant, the differential of $\Phi_{g}$ maps vertical tangent spaces with respect to $p$ isomorphically onto one another for each $g\in G$. Thus,
@@
$\left<\nabla f(x),v\right>=\left<\nabla f(h(x)),Dh_{x}(v)\right>\ \ \forall x\in E|_{U},\ v\in T_{x}E.$
@@
$\left<\nabla f(x),v\right>_{x}$ $=\left<(D\Phi_{\eta(x)})_{x}(\nabla^{\mathrm{ver}}f(x)),(D\Phi^{x})_{\eta(x)}(D\eta_{x}(v))+(D\Phi_{\eta(x)})_{x}(v)\right>_{h(x)}$
$=\left<\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(v)+v\right>_{x},$
@@
$\|\nabla f(x)\|^{2}$ $=\left<\nabla^{\mathrm{ver}}f(x),(D\Phi_{\eta(x)})_{x}^{-1}(D\Phi^{x})_{\eta(x)}D\eta_{x}(\nabla f(x))\right>_{x}+\|\nabla^{\mathrm{ver}}f(x)\|^{2}$
$\leq\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|(D\Phi^{x})_{\eta(x)}\|_{\mathrm{op}}\cdot\|D\eta_{x}\|_{\mathrm{op}}\cdot\|\nabla f(x))\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}.$
where have used the Cauchy-Schwarz inequality and the $G$-invariance of the metric. Since $\Phi$ is $f$-controlled with respect to $C$, we obtain
$\|\nabla f(x)\|^{2}$ $\leq C(f(x))\cdot\|\nabla^{\mathrm{ver}}f(x)\|\cdot\|D\eta_{x}\|_{\mathrm{op}}\cdot\|\nabla f(x)\|+\|\nabla^{\mathrm{ver}}f(x)\|^{2}.$
@@
Since the chosen metric on $G$ is bi-invariant, this yields that $\|D\mathrm{inv}_{g}\|_{\mathrm{op}}=1$ for all $g\in G$. We derive that $\|D\eta_{x}\|_{\mathrm{op}}\leq\|Ds_{p(x)}\|_{\mathrm{op}}$ for all $x\in f^{\lambda}\cap E|_{U}$. Inserting this into the above inequality shows the claim. ∎
@@
[…]
@@
[…]
In the rest of this section, we want to apply Corollary 8.8 to unit tangent bundles of odd-dimensional spheres. The parametrized TC of unit tangent bundles of odd-dimensional spheres has been computed by Minowa in [Min25]. A general upper bound of the parametrized TCs of unit sphere bundles of vector bundles has been obtained by Farber and Weinberger in [FW23] and extended to sequential TCs by Farber and Paul in [FP24]. Using our approach, we will compute the values of the sequential parametrized TCs of unit tangent bundles of (4m1)(4m-1)-dimensional spheres, where $m\in\mathbb{N}$.
In the following we fix $n\in\mathbb{N}$ and will use the canonical identification $\mathbb{R}^{2n}\cong\mathbb{C}^{n}$ without always spelling it out explicitly. We further view $\mathbb{S}^{2n-1}$ as equipped with the metric induced by the standard metric on $\mathbb{R}^{2n}$ and denote its unit tangent bundle by $p\colon U\mathbb{S}^{2n-1}\to\mathbb{S}^{2n-1}.$ Explicitly,
@@
where $\left$$\cdot$$\cdot\right$> denotes the Euclidean inner product on $\mathbb{R}^{2n}$. Seen as a subgroup of $SO(2n)$, the special unitary group $SU(n)$ acts transitively on the sphere. It further induces an action on $U\mathbb{S}^{2n-1}$, which is not transitive. The map $p$ is $SU(n)$-equivariant and scalar multiplication with the complex unit induces an $SU(n)$-equivariant section
@@
(8.5) x1,y1=x2,y2=0as well as $\langle x_{1},y_{1}\rangle=\langle x_{2},y_{2}\rangle=0\quad\text${as well as $\quad\langle$ $\rangle+\langle x_{1},y_{2}\rangle$=0.
@@
(8.7) $\operatorname{\mathsf{TC}}_{r}[p]\leq\sum_{\mu\in\mathbb{R}}\max\Big\{\operatorname{\mathsf{TC}}_{r,p}(\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{2n-1}}\dots\times_{\mathbb{S}^{2n-1}}\times\operatorname{Crit}_{\mu_{r}}f)\ \Big|\ \sum_{j=1}^{r}\mu_{j}=\mu\Big\},$
where $\operatorname{\mathsf{TC}}_{r}[p]:=\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{2n-1}\to\mathbb{S}^{2n-1}]$. We want to further compute this upper bound of $TC_{r}[p]$ and study the critical sets of $f$ in greater detail for this purpose.
@@
To show the converse inclusion, let $X=(x_{1},x_{2})\in U\mathbb{S}^{2n-1}$ with $x_{2}\notin\{ix_{1},-ix_{1}\}$. Set $Y=(0,w)$ with $w=ix_{1}-\langle ix_{1},x_{2}\rangle x_{2}$. One verifies that indeed $Y\in T_{X}U\mathbb{S}^{2n-1}$ and it holds that $Df_{X}(Y)=1-\langle ix_{1},x_{2}\rangle^{2}\neq 0$. Hence, $X\not\in\operatorname{Crit}f$, which shows that any critical point of $f$ is of the form derived above. ∎
Let $r\in\mathbb{N}$ with $r\geq 2$. We have seen in the previous Lemma that the critical values of $f$ are $1$ and 1-1, so the fiber products of the critical manifolds occurring in (8.7) are of the form
@@
We want to explicitly compute the sequential parametrized subspace complexities that occur on the right-hand side of this inequality in the case of even $n$, i.e. for spheres of dimension 4m14m-1 for some $m\in\mathbb{N}$. Eventually, we will show the following.
@@
For each $m\in\mathbb{N}$, it holds that $\operatorname{\mathsf{TC}}_{r}[p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}]=r+1.$
@@
In particular, it holds that $\langle x,jx\rangle=\langle ix,jx\rangle=0$ and $\|jx\|^{2}=\|x\|^{2}$ for all $x\in\mathbb{R}^{4m}$.
@@
Let $m\in\mathbb{N}$ and let $f\colon\mathbb{S}^{4m-1}\to\mathbb{R}$ be given as in (8.6). Let $\mu_{1},\ldots,\mu_{r}\in\{+1,-1\}$ and consider the manifold $\mathcal{C}=\operatorname{Crit}_{\mu_{1}}f\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{2}}f\times_{\mathbb{S}^{4m-1}}\ldots\times_{\mathbb{S}^{4m-1}}\operatorname{Crit}_{\mu_{r}}f.$ Then $\operatorname{\mathsf{TC}}_{r,p}(\mathcal{C})=1$.
@@
Consider the map $B\colon\mathbb{S}^{4m-1}\to U\mathbb{S}^{4m-1}$, $B(x)=jx$. The map $B$ satisfies $\langle A(x),B(x)\rangle=0$ for all $x\in\mathbb{S}^{4m-1}$. Let $u=((x,v_{1}),\ldots,(x,v_{r}))\in\mathcal{C}$, so that $v_{j}\in\{ix,-ix\}$ for all $j\in\{1,2,\ldots,r\}$. We define a path $\sigma_{u}\colon I\to U\mathbb{S}^{n}$ as follows. For all $j\in\{0,1,\ldots,r-2\}$ and $t\in[\frac{j}{r-1},\frac{j+1}{r-1}]$ we put
@@
The fibration $p\colon U\mathbb{S}^{4m-1}\to\mathbb{S}^{4m-1}$ is fiber homotopy equivalent to a trivial fibration if and only if $m\in\{1,2\}$, so that Theorem 8.10 can not be derived by elementary arguments for $m\geq 3$. More precisely, as shown in [JW54, Theorem 1.1], $U\mathbb{S}^{4m-1}$ has the homotopy type of $\mathbb{S}^{4m-1}\times\mathbb{S}^{4m-2}$ if and only if there exists an element of Hopf invariant one in $\pi_{8m-1}(\mathbb{S}^{4m})$, which By [Ada60] this is the case only for $m\in\{1,2\}$. Thus, $p$ is not fiber homotopy trivial for $m\geq 3$.